What is the probability that the top 10 cards in a shuffled deck are red? What is the probability that the top k cards in a shuffled deck are red?
Answer #1

probabiltiy that top 10 cards are red =P(10 cards 10 from 13 and 0 from rest of 39)

=13C10 /52C10   =(13!/(3!*10!))/(52!/(10!*42!)) =(13!*42!)/(52!*3!)=0.00000001808

probability that the top k cards in a shuffled deck are red =13Ck /52Ck   where k<=13